OpenStudy (anonymous):

Which of the following is an incorrect expression of a trig function in terms of sin 0 (lines through all zeros) a. csc 0 = 1 / sin 0 b. cot 0 = sqrt. 1-sin^20 / sin 0 c. tan 0 = sin 0 / sqrt. 1-sin ^2 0 d. sec 0 = 1/ sqrt. 1-sin^2 0

OpenStudy (anonymous):

@aum :)

OpenStudy (aum):

csc(x) = 1 / sin(x). Therefore, a is correct. (Note: They want the incorrect choice). Can you try the other choices?

OpenStudy (anonymous):

How do I do that? Do I use my calculator?

OpenStudy (aum):

No, you use trig identities.

OpenStudy (aum):

sqrt( 1 - sin^20) / sin 0 ) sqrt( 1 - sin^2(x)) / sin(x) ) = sqrt( cos^2(x) ) / sin(x) = cos(x) / sin(x) = cot(x). Therefore, b is correct.

OpenStudy (aum):

Did you try choices c and d. Remember. they want the choice that is incorrect. So a and b are not the answer because they are correct.